OpenStudy (anonymous):

(x2 + 13x + 40) ÷ (x + 8) show steps please

OpenStudy (anonymous):

does it help to know that \[x^2+13x+40=(x+8)(x+5)\]?

OpenStudy (anonymous):

no..could you work it out please?

OpenStudy (anonymous):

since the numerator factors as \[(x+8)(x+5)\] then \[\frac{x^2+13x+40}{x+8}=\frac{(x+8)(x+5)}{x+8}= \frac{\cancel{(x+8)}(x+5)}{\cancel{x+8}}=x+5\]

OpenStudy (anonymous):

Okay, I have a better understanding now. thanks

OpenStudy (anonymous):

yw